OpenStudy (anonymous):

How does t2−15 Becomes (t−15−−√)(t+15−−√)

OpenStudy (anonymous):

hey w8 ill write that again

OpenStudy (anonymous):

how does \[t ^{2}-15\] becomes \[(t-\sqrt{15})(t+\sqrt{15})\]

OpenStudy (amistre64):

it doesnt "become" it

OpenStudy (anonymous):

\[(a+b)(a-b)=a^2+ba-ba-b^2\]\[=a^2-b^2\] \[b^2=15 \therefore b=\sqrt{15}\] just substitute

OpenStudy (anonymous):

i didnt get it

OpenStudy (unklerhaukus):

± @Blueskys

OpenStudy (anonymous):

@UnkleRhaukus sorry about that

OpenStudy (anonymous):

I got it thanx